Engineering Mangement and Entrepreneurship

Find equilibrium Price and Quantity and Price Elasticity of Demand from the

following demand and supply functions

i. Q=10-2P 2  Q=5+3P

ii. Qd= 200-40P Qs= -40+80P
Answer #1

ANSWER :


(I)


Demand function :  Q = 10 - 2P


Supply function : Q = 5 + 3P 


At equilibrium, Demand quantity = Supply quantity 


So,


10 - 2P = 5 + 3P 

=> 10 - 5 = 3P + 2P

=> 5 = 5P

=> P = 1 (currency unit) 

=> Q = 10 - 2*1 = 8 units 


So, equilibrium Price is 1 (currency unit) and equilibrium Quantity is 8 units.

(ANSWER).


(Ii)


Demand function :  Qd = 200 - 40P


Supply function : Qs = - 40 + 80P


At equilibrium, Qd = Qs


So,


200 - 40P = - 40 + 80P

=> 200 + 40 = 80P + 40P

=> 240 = 120P

=> P = 2 (currency unit) 

=> Q = 200 - 40(2) = 120 units


So, equilibrium Price is 2 (currency unit) and equilibrium Quantity is 120 units.

(ANSWER).


answered by: Tulsiram Garg